write an exponential function of the form y=ab^x whos graph passes trough the points, (2,5.888),(3,9.4208)
Answer by stanbon(48569)

write an exponential function of the form y=ab^x whos graph passes trough the points, (2,5.888),(3,9.4208)
--------------
5.888 = ab^2
9.4208 = ab^3
------------
Divide the 1st equation into the 2nd to get:
1.6 = b
Now, solve for "a" by substituting 1.6 for b
5.888 = a (1.6)^2
a= 5.888/(1.6)^2 = 2.3
---------
So, y=(2.3)(1.6)^x
--------
Cheers,
Stan H.